find "y'" for "y= x\u00b3 (sqrt(x\u00b2 + 1))"


1
Expert's answer
2021-02-02T14:59:38-0500

"y=x^3\cdot\sqrt{x²+1}"


"y'=(x^3)'\cdot\sqrt{x^2+1}+x^3\cdot(\sqrt{x^2+1})'="


"\displaystyle 3x^2\cdot\sqrt{x^2+1}+x^3\cdot\frac{1}{2\sqrt{x^2+1}}\cdot(x^2+1)'="


"\displaystyle 3x^2\sqrt{x^2+1}+x^3\cdot\frac{1}{2\sqrt{x^2+1}}\cdot2x="

"\displaystyle 3x^2\sqrt{x^2+1}+\frac{x^4}{\sqrt{x^2+1}}="


"\displaystyle 3x^2\sqrt{x^2+1}\cdot\frac{\sqrt{x^2+1}}{\sqrt{x^2+1}}+\frac{x^4}{\sqrt{x^2+1}}="


"\displaystyle \frac{3x^2(x^2+1)}{\sqrt{x^2+1}}+\frac{x^4}{\sqrt{x^2+1}}="


"\displaystyle \frac{3x^4+3x^2+x^4}{\sqrt{x^2+1}}="


"\displaystyle \frac{4x^4+3x^2}{\sqrt{x^2+1}}"